# From Introspection to Best Practices: Principled Analysis of Demonstrations in Multimodal In-Context Learning

### v1

@@ Abstract
Motivated by in-context learning (ICL) capabilities of Large Language models (LLMs), multimodal LLMs with additional visual modality are also exhibited with similar ICL abilities when multiple image-text pairs are provided as demonstrations. However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works. We conduct a systematic and principled evaluation of multimodal ICL for models of different scales on a broad spectrum of new yet critical tasks. Through perturbations over different modality information, we show that modalities matter differently across tasks in multimodal ICL. Considering such modality impact, we further utilize modality-driven demonstration strategies to boost ICL performance. We also identify that demonstration selection is closely related to the models’ ability to capture task inductive biases from multimodal ICL. Our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal in-context learning, and sheds light on effectively improving multimodal ICL on a wide range of tasks even if those tasks are not seen in or even contradict pretraining data.
@@ 1 Introduction
Figure 1: (b) Modality matters differently on ICL across tasks: visual information matters little on Application but a lot on KIE, textual answers are more important to ICL on KIE than that to Application. (c) Demonstrations selected by text-driven strategy BERTScore benefits more on Application, while those selected by visual similarity (CLIP) bring higher accuracy on KIE.
@@
Motivated by in-context learning (ICL) capabilities of Large Language Models (LLMs) for NLP tasks (Brown et al., 2020; Garg et al., 2022; Akyürek et al., 2022), multimodal LLMs with additional visual modality are also exhibited with similar ICL abilities when multiple image-text pairs are provided as demonstrations (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). In recent studies, the Retrieval-based In-Context Example Selection (RICES, Yang et al. (2022)) approach, which retrieves similar images in the support set by comparing their visual features with testing images, has become a default approach to select demonstrations for multimodal in-context learning (Alayrac et al., 2022; Sun et al., 2023; Yang et al., 2024).
@@
However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works, nor there has been enough justification for the necessity of selecting demonstrations according to visual modality and analyze its advantages over other modalities. Yang et al. (2024) only explored better in-context configurations for image captioning, while Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations. However, their observations are limited to image captioning (Young et al., 2014; Chen et al., 2015) and general-purpose visual question answering tasks (Goyal et al., 2017; Gurari et al., 2018; Marino et al., 2019; Sidorov et al., 2020), which leaves a comprehensive exploration for the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs.
@@
In this paper, we conduct a systematic and principled evaluation of multimodal ICL for models of different scales (ranging from OpenFlamingo 4B, Awadalla et al. (2023)) to IDEFICS1 80B, Laurençon et al. (2023)) on a broad spectrum of new yet critical tasks as shown in Figure 1(a). These tasks require different types of capabilities, including hallucination mitigation (Wang et al., 2023a), text-rich image understanding (Liu et al., 2023; Li et al., 2024), medical information comprehension (He et al., 2020; Pacheco et al., 2020; Liu et al., 2021), and cross-style transfer (Cai et al., 2023), etc. With diverse ICL capabilities examination, we show that the dependency of performance gain from ICL on demonstration modalities differs among tasks (Section 4). As demonstrated in Figure 1(b), perturbing visual information in demonstrations (e.g., removing or replacing with random, noised or permuted images) does not cause significant performance drop on ICL for cross-style and medical tasks, while resulting in decreased accuracy than that provided by correct demonstrations on tasks such as key information extraction from text-rich images. Sometimes it even leads to much worse performance than the zero-shot inference. On the other hand, textual perturbations (e.g., replacing the question/answer with random or one from other candidates in the same demonstration set) hurt ICL performance to different extents across tasks. For instance, perturbations on either questions or answers lead to greatly reduced accuracy on some tasks, while perturbations on answers results in extremely bad performance on others. These observations strongly suggest the necessity of understanding modality impact on ICL prior to collecting demonstrations for specific tasks.
@@
We conduct further investigation on how to select effective demonstration to boost multimodal ICL performance (Section 5). As shown in Figure 1(c), we identify that providing demonstrations selected by textual similarity (e.g., text encoder of CLIP (Radford et al., 2021) or BERTScore (Zhang et al., 2019)) benefits ICL performance consistently across models and tasks. This is consistent with literature (Chen et al., 2023) and our prior observation that the textual modality plays an important role in ICL performance. For tasks observed with vital impact from visual modality on ICL performance, demonstrations selected by visual similarity (e.g., vision encoder of CLIP) elicit drastically improved ICL performance. Moreover, demonstration selection strategies that consider both visual and textual modalities, such as ALBEF (Li et al., 2021) with a multimodal encoder that explicitly models interactions between image and text features, present trade-off performance regardless of various modality importance to specific tasks.
@@
Lastly, we illustrate that models may not always capture task inductive biases from multimodal ICL (Section 6). Concretely, we flip annotations of demonstrations to override semantic priors learned during pretraining (e.g., “Yes” to admit hallucinated objects in images and “No” to deny the presence actually existing objects in images). Small-scale models fail to comprehend or follow practices against prior knowledge provided by randomly sampled demonstrations. Surprisingly, models learn to follow inductive biases given demonstrations selected according to textual similarities, an emergent ability unlocked by scaling studied in literature (Wei et al., 2022b; Zhou et al., 2022). This is reasonable as flipped annotations mainly convey inductive biases through texts. Such capability of capturing inductive bias of demonstrations without scaling up models is more attractive than using semantic priors, since the model would be able to perform a wide range of tasks without further tuning, even if those tasks are not seen in or even contradict pretraining data (Wei et al., 2023).
@@
In summary, our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal in-context learning. We empirically show that (1) modalities matter differently in multimodal ICL across tasks (Section 4), (2) demonstration strategies considering modality impact are able to boost ICL performance (Section 5), (3) demonstration selection is closely related to models’ ability to capture task inductive biases from multimodal ICL (Section 6). Overall, our work aims to shed light on effectively improving multimodal ICL on a wide range of tasks even if those tasks are not seen in or even contradict pretraining data.
@@ 2 Related Work
LLMs have been recognized as strong few-shot learners since their emergence (Brown et al., 2020). With ICL, LLMs are empowered to generalize to a wide range of tasks at inference even if those tasks are not seen in pretraining data (Garg et al., 2022; Akyürek et al., 2022). However, the performance of ICL is critically sensitive to the choices of demonstrations (Rubin et al., 2022; Wang et al., 2023b; Gupta et al., 2023), the order (Lu et al., 2022; Wu et al., 2023) and format of prompts (Zhao et al., 2021; Min et al., 2021).
To understand why ICL works, Xie et al. (2021) explained from the theoretical perspective that Transformers acquire the ICL ability when they are trained to infer latent concepts during pretraining. Min et al. (2022) empirically showed that the performance gain of ICL over zero-shot inference is mainly driven by the label space, distribution of input text, output labels, and overall format of the sequence, while the represented mapping from inputs to the outputs in demonstrations matters little. However, some recent work (Zhou et al., 2022; Wei et al., 2023) suggested that when scaling up to some extent, larger models (e.g. PaLM-540B (Chowdhery et al., 2023) and Codex (Chen et al., 2021)) can actually learn input-output mappings, which allows them to perform a variety of challenging tasks even if they contradict pretraining data.
@@
After pretraining on interleaved image-text data or fine-tuning on multi-turn conversations, multimodal LLMs have exhibited ICL abilities in tasks such as image captioning and general-purpose visual question answering (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). Considering these studies may not sufficiently reveal strengths and weaknesses of ICL, Zong et al. (2024) recently introduced VL-ICL Bench which encompasses a broad spectrum of tasks for multimodal ICL evaluation. However, there is not much work that conducts principled analysis on emergent ICL capabilities and provides insightful suggestions for future ICL practices. Yang et al. (2024) only explored better in-context configurations for image captioning.
One work that is closely connected to ours is Chen et al. (2023). Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations and proposed Mixed Modality In-Context Example Selection (MMICES), which first pre-filters samples based on visual feature similarity and then selects most similar ones based on textual similarity. However, their observations are limited to image captioning and general-purpose visual question answering tasks, which leaves a comprehensive exploration for the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs. We conduct more comprehensive study on the impact of modality to ICL and find that modalities matter differently across tasks. Furthermore, we investigate how models of different scales capture task inductive biases from multimodal ICL.

### v3

@@ Abstract
Motivated by in-context learning (ICL) capabilities of Large Language Models (LLMs), multimodal LLMs with additional visual modality exhibit similar ICL abilities when multiple image-text pairs are provided as demonstrations. However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works. We conduct a systematic and principled evaluation of multimodal ICL for models of different scales on a broad spectrum of new yet critical tasks. Through perturbations over different modality information, we show that modalities matter differently across tasks in multimodal ICL. Guided by task-specific modality impact, we recommend modality-driven demonstration strategies to boost ICL performance. We also find that models may follow inductive biases from multimodal ICL even if they are rarely seen in or contradict semantic priors from pretraining data. Our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal in-context learning, and sheds light on effectively improving multimodal ICL on a wide range of tasks.  [^1] Codes and datasets are available at https://github.com/luka-group/MultimodalICLBestPractice.
@@ 1 Introduction
Figure 1: (b) Modality matters differently on ICL across tasks: visual information matters little on Application but a lot on KIE, textual answers are more important to ICL on KIE than that on Application. (c) Demonstrations selected by text-driven strategy BERTScore benefit more on Application, while those selected by visual similarity (CLIP) bring higher accuracy on KIE.
@@
Motivated by in-context learning (ICL) capabilities of Large Language Models (LLMs) for NLP tasks (Brown et al., 2020; Garg et al., 2022; Akyürek et al., 2022), multimodal LLMs with additional visual modality are also exhibited with similar ICL abilities when multiple image-text pairs are provided as demonstrations (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). In recent studies, the Retrieval-based In-Context Example Selection (RICES, Yang et al. (2022)) approach, which retrieves similar images in the support set by comparing their visual features with testing images, has become a default approach to selecting demonstrations for multimodal ICL (Alayrac et al., 2022; Sun et al., 2023; Yang et al., 2024).
Figure 2: Benchmarks with inductive bias contracting the semantic priors (left) or rarely seen in pretraining data (middle and right). We list ground-truth (GT), zero-shot responses from GPT-4o and provide ICL analysis in Section 6.
@@
However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works, nor has there been enough justification for the necessity of selecting demonstrations according to visual modality and analyzing its advantages over other modalities. Yang et al. (2024) only explored better in-context configurations for image captioning, while Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations. However, their observations are limited to image captioning (Young et al., 2014; Chen et al., 2015) and general-purpose visual question answering tasks (Goyal et al., 2017; Gurari et al., 2018; Marino et al., 2019; Sidorov et al., 2020), which leaves a comprehensive exploration of the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs.
In this paper, we conduct a systematic and principled evaluation of multimodal ICL for models of different scales (ranging from OpenFlamingo 4B, Awadalla et al. (2023) to IDEFICS1 80B, Laurençon et al. (2023)) on a broad spectrum of new yet critical tasks as shown in Figure 1(a). These tasks require different types of capabilities, including hallucination mitigation (Wang et al., 2023), text-rich image understanding (Liu et al., 2023; Li et al., 2024), medical information comprehension (He et al., 2020; Pacheco et al., 2020; Liu et al., 2021), and cross-style transfer (Cai et al., 2023), etc.
@@
With diverse ICL capabilities examination, we show that the dependency of performance gain from ICL on demonstration modalities differs among tasks (Section 4). For example, perturbing visual information in demonstrations (e.g., removing or replacing with random, noised or permuted images) does not cause significant performance drop on ICL for Application task (Figure 1(b) left), while resulting in decreased accuracy than that provided by correct demonstrations on tasks such as key information extraction (KIE) from text-rich images (Figure 1(b) right). On the other hand, textual perturbations (e.g., replacing the question/answer with random or one from other candidates in the same demonstration set) hurt ICL performance to different extents across tasks. These observations strongly suggest the necessity of understanding modality impact on ICL prior to collecting demonstrations for specific tasks.
@@
We conduct further investigation on how to select effective demonstrations to boost multimodal ICL performance (Section 5). Based on empirical experiments, we recommend the following practices to elicit better ICL performance from multimodal models: 1) Utilizing demonstrations selected by visual similarity (e.g., vision encoder of CLIP) for tasks observed with vital impact from visual modality on ICL performance, e.g., KIE task shown on the right of Figure 1(c). 2) Selecting demonstrations with high textual similarity (e.g., text encoder of CLIP (Radford et al., 2021) or BERTScore (Zhang et al., 2019)) for tasks if the textual modality plays an important role in ICL performance, e.g., Application task shown on the left of Figure 1(c). 3) Choosing demonstrations with both visual and textual similarity considered (e.g., ALBEF (Li et al., 2021) with a multimodal encoder that explicitly models interactions between image and text features) if dual modalities matter similarly to ICL performance.
@@
Lastly, we illustrate that models may have the capability to capture task inductive biases from multimodal ICL (Section 6). We investigate two kinds of inductive bias: 1) contradicting semantic priors, and 2) rarely seen in pretraining. By deliberately flipping annotations of demonstrations to override strong semantic priors learned during pretraining ( Figure 2 left), small-scale models fail to comprehend or follow practices provided by randomly sampled demonstrations, while they learn to follow inductive biases given demonstrations selected according to textual similarities, an emergent ability unlocked by scaling studied in literature (Wei et al., 2022b; Zhou et al., 2022). Although multimodal LLMs are typically pretrained on true positives only but rarely on hallucination-inducing scenarios with unanswerable questions ( Figure 2 middle,  Cha et al. (2024)), multimodal ICL greatly reduces hallucination over zero-shot inference. We also observe effectiveness of multimodal ICL in addressing the failure mode (Ball et al., 2024; Shin and Kaneko, 2024; Yehudai et al., 2024) of models on tasks humans find trivial (e.g., counting r’s in the image plotting “strawberry” as shown in  Figure 2 right). Such capability to capture inductive bias of demonstrations without scaling up models is more attractive than using semantic priors, since the model would be able to perform a wide range of tasks without further tuning, even if those tasks are not seen in or even contradict pretraining data.
@@
In summary, our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal ICL. We empirically show that (1) modalities matter differently in multimodal ICL across tasks (Section 4), (2) demonstration strategies considering modality impact are able to boost ICL performance (Section 5), (3) models are capable of capturing task inductive biases from multimodal ICL (Section 6). Overall, our work aims to shed light on effectively improving multimodal ICL on a wide range of tasks even if those tasks are not seen in or even contradict pretraining data.
@@ 2 Related Work
LLMs have been recognized as strong few-shot learners since their emergence (Brown et al., 2020). With ICL, LLMs are empowered to generalize to a wide range of tasks at inference even if those tasks are not seen in pretraining data (Garg et al., 2022; Akyürek et al., 2022). To understand why ICL works, Min et al. (2022) empirically showed that the performance gain of ICL over zero-shot inference is mainly driven by the label space, distribution of input text, output labels, and overall format of the sequence, while the represented mapping from inputs to the outputs in demonstrations matters little. However, some recent work (Zhou et al., 2022; Wei et al., 2023) suggested that when scaling up to some extent, larger models can actually learn input-output mappings, which allows them to perform a variety of challenging tasks even if they contradict pretraining data.
@@
After pretraining on interleaved image-text data or fine-tuning on multi-turn conversations, multimodal LLMs have exhibited ICL abilities in tasks such as image captioning and general-purpose visual question answering (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). Considering these studies may not sufficiently reveal strengths and weaknesses of ICL, Zong et al. (2024) recently introduced VL-ICL Bench which encompasses a broad spectrum of tasks for multimodal ICL evaluation. However, there is not much work that conducts principled analysis on emergent ICL capabilities and provides insightful suggestions for future ICL practices. Yang et al. (2024) only explored better in-context configurations for image captioning. Qin et al. (2024) recognized three factors–demonstration retrieval, ordering and instructions, that contribute to multimodal ICL performance, without providing task-specific suggestions for demonstration selection to boost multimodal ICL capabilities.
One work that is closely connected to ours is Chen et al. (2023). Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations and proposed Mixed Modality In-Context Example Selection (MMICES), which first pre-filters samples based on visual feature similarity and then selects most similar ones based on textual similarity. However, their observations are limited to image captioning and general-purpose visual question answering tasks, which leaves a comprehensive exploration for the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs. We conduct more comprehensive study on the impact of modality on ICL and find that modalities matter differently across tasks. Furthermore, we investigate how models of different scales capture task inductive biases from multimodal ICL.